OpenStudy (anonymous):

solve using complete the square: 2x^2-4x-8=0

jimthompson5910 (jim_thompson5910):

Hint: 2x^2-4x-8=0 2x^2-4x=8 2(x^2-2x)=8 2(x^2-2x+1-1)=8 2((x^2-2x+1)-1)=8 2(x^2-2x+1)-2=8 2(x^2-2x+1)=8+2 2(x^2-2x+1)=10 2(x-1)^2=10

OpenStudy (anonymous):

why did you do +1 and -1

jimthompson5910 (jim_thompson5910):

because the +1-1 is really +0

jimthompson5910 (jim_thompson5910):

and that doesn't change anything

jimthompson5910 (jim_thompson5910):

you add and subtract 1 to complete the square

OpenStudy (mathstudent55):

@jim_thompson5910 Why didn't you divide both sides by 2 as your first step?

jimthompson5910 (jim_thompson5910):

you can do that, but i just kept it for some reason it doesn't matter either way

OpenStudy (mathstudent55):

Ok, thanks. I was just wondering if there was a specific reason for not doing it.

jimthompson5910 (jim_thompson5910):

no i guess i just forgot about it and copied/pasted it down the line